# An LLM-Powered Agent for Physiological Data Analysis: A Case Study on PPG-based Heart Rate Estimation

### v1

@@ Abstract
Large language models (LLMs) are revolutionizing healthcare by improving diagnosis, patient care, and decision support through interactive communication. More recently, they have been applied to analyzing physiological time-series like wearable data for health insight extraction. Existing methods embed raw numerical sequences directly into prompts, which exceeds token limits and increases computational costs. Additionally, some studies integrated features extracted from time-series in textual prompts or applied multimodal approaches. However, these methods often produce generic and unreliable outputs due to LLMs’ limited analytical rigor and inefficiency in interpreting continuous waveforms. In this paper, we develop an LLM-powered agent for physiological time-series analysis aimed to bridge the gap in integrating LLMs with well-established analytical tools. Built on the OpenCHA, an open-source LLM-powered framework, our agent features an orchestrator that integrates user interaction, data sources, and analytical tools to generate accurate health insights. To evaluate its effectiveness, we implement a case study on heart rate (HR) estimation from Photoplethysmogram (PPG) signals using a dataset of PPG and Electrocardiogram (ECG) recordings in a remote health monitoring study. The agent’s performance is benchmarked against OpenAI GPT-4o-mini and GPT-4o, with ECG serving as the gold standard for HR estimation. Results demonstrate that our agent significantly outperforms benchmark models by achieving lower error rates and more reliable HR estimations. The agent implementation is publicly available on GitHub [^1] https://github.com/mohammadfeli/CHA-PPGHR.
@@ I INTRODUCTION
Large language models (LLMs) have emerged as transformative tools in healthcare, fundamentally reshaping the interaction between medical professionals, patients, and medical data [1]. Their ability to understand and generate human-like text has led to widespread adoption across various applications, including medical diagnosis, patient care, clinical decision support, drug discovery, health recommendation systems, and healthcare chatbots [2]. By enabling personalized and efficient communication, LLMs provide more interactive and proactive healthcare solutions. They are typically implemented using prompt engineering, which involves crafting precise input prompts to refine responses, and fine-tuning, which customizes the model for specific healthcare applications [3].
@@
To integrate LLMs with physiological time-series data, three primary strategies have been explored in the literature. The first approach involves directly incorporating numerical time-series sensor data into textual prompts to perform various analytical tasks using LLMs [5, 6, 4, 7, 8]. However, these methods struggle with handling large numerical sequences, often exceeding token limits in LLMs and increasing computational costs. The second approach extracts features from time-series health data and integrates them into textual prompts [9, 10, 11, 12]. While this reduces token overhead, it still leads to generic and unreliable outputs, due to LLMs’ limitations in analytical reasoning. The third approach, multimodal prompting, introduces visual representations of physiological signals, such as Electrocardiogram (ECG) and Photoplethysmogram (PPG), for LLM-based biosignal analysis [13, 14]. Similar to feature-based prompting, these methods also fail to provide analytically grounded insights, as LLMs are not inherently designed to interpret continuous waveforms directly.
@@ II Related Works
Studies utilizing textual prompting with direct integration of time-series data represent the numerical data in a structured text format for the model. One of the earliest studies in this context [5] demonstrated the use of LLMs for tasks such as arrhythmia detection and physical activity recognition by directly integrating interbeat interval sequences and accelerometer sensor data into prompts. Similarly, Kim et al. [6] explored LLMs for health prediction by incorporating various sensor readings into queries, such as step count, calories burned, and resting heart rate. Chan et al. [4] proposed an LLM-based framework for tasks such as semantic segmentation, boundary detection, and anomaly detection, embedding biosignals directly into model inputs. Additionally, Imran et al. [7] employed LLMs for physical activity analysis, where gyroscope and accelerometer sensor data were incorporated into prompts. LLMs have also been applied to cardiac activity monitoring, metabolic health prediction, and sleep detection by leveraging wearable sensor data [8]. However, LLMs are not inherently specialized for processing long raw numerical sequences. Physiological time-series data often consist of thousands or even millions of data points, and representing them as textual input for LLMs can result in poor performance. This approach also exceeds the token limits of most LLMs and leads to high computational costs due to the large number of tokens.
@@ IV Evaluation and Results
We evaluate our LLM-powered agent using the dataset described in Section III-B1. For this evaluation, we use PPG data from 12 participants, while data from the remaining 34 participants was previously utilized in the development of machine learning models within the PPG processing pipeline. In total, the evaluation dataset consists of 621 PPG recordings, each spanning 15 minutes.
@@ IV-B Gold reference and comparison criteria
For evaluation, we compute Mean Absolute Error (${MAE}=\frac{1}{n}\sum_{i=1}^{n}|y_{i}-\hat{y}_{i}|$), Root Mean Square Error (${RMSE}=\sqrt{\frac{1}{n}\sum_{i=1}^{n}(y_{i}-\hat{y}_{i})^{2}}$), Mean Absolute Percentage Error (${MAPE}=\frac{100\%}{n}\sum_{i=1}^{n}\left|\frac{y_{i}-\hat{y}_{i}}{y_{i}}\right|$), and Median Absolute Difference (${MAD}=\text{median}(|y_{i}-\hat{y}_{i}|)$), where $n$ is the number of samples, $y_{i}$ represents the ground-truth values, and $\hat{y}_{i}$ is the predicted value of HR or. Additionally, we perform linear regression analysis to assess the linear relationship between estimated and reference HR values and use the Bland-Altman method to evaluate the agreement between the estimated and ground truth HR values.

### v2

@@ Abstract
Large language models (LLMs) are revolutionizing healthcare by improving diagnosis, patient care, and decision support through interactive communication. More recently, they have been applied to analyzing physiological time-series like wearable data for health insight extraction. Existing methods embed raw numerical sequences directly into prompts, which exceeds token limits and increases computational costs. Additionally, some studies integrated features extracted from time-series in textual prompts or applied multimodal approaches. However, these methods often produce generic and unreliable outputs due to LLMs’ limited analytical rigor and inefficiency in interpreting continuous waveforms. In this paper, we develop an LLM-powered agent for physiological time-series analysis aimed to bridge the gap in integrating LLMs with well-established analytical tools. Built on the OpenCHA, an open-source LLM-powered framework, our agent powered by OpenAI’s GPT-3.5-turbo model features an orchestrator that integrates user interaction, data sources, and analytical tools to generate accurate health insights. To evaluate its effectiveness, we implement a case study on heart rate (HR) estimation from Photoplethysmogram (PPG) signals using a dataset of PPG and Electrocardiogram (ECG) recordings in a remote health monitoring study. The agent’s performance is benchmarked against OpenAI GPT-4o-mini and GPT-4o, with ECG serving as the gold standard for HR estimation. Results demonstrate that our agent significantly outperforms benchmark models by achieving lower error rates and more reliable HR estimations. The agent implementation is publicly available on GitHub [^1] https://github.com/mohammadfeli/CHA-PPGHR.
@@ I INTRODUCTION
Large language models (LLMs) have emerged as transformative tools in healthcare, reshaping the interaction between medical professionals, patients, and medical data [1]. Their ability to understand and generate human-like text has led to widespread adoption across various applications, including medical diagnosis, patient care, clinical decision support, drug discovery, health recommendation systems, and healthcare chatbots [2]. By enabling personalized and efficient communication, LLMs provide more interactive and proactive healthcare solutions. They are typically adapted using prompt engineering, which involves crafting precise input prompts to refine responses, and fine-tuning, which customizes the model for specific healthcare applications [3].
@@
To integrate LLMs with physiological time-series data, three primary strategies have been explored in the literature. The first approach involves directly incorporating numerical time-series sensor data into textual prompts to perform various analytical tasks using LLMs [5, 6, 4, 7, 8]. However, these methods struggle with handling large numerical sequences, often exceeding token limits in LLMs and increasing computational costs. The second approach extracts features from time-series health data and integrates them into textual prompts [9, 10, 11, 12]. While this reduces token overhead, it still leads to generic and unreliable outputs, due to LLMs’ limitations in analytical reasoning. The third approach, multimodal prompting, introduces visual representations of physiological signals, such as Electrocardiogram (ECG) and Photoplethysmogram (PPG), for LLM-based biosignal analysis [13, 14]. Similar to feature-based prompting, these methods also fail to provide analytically grounded insights, as LLMs are not inherently designed to interpret continuous waveforms.
@@ II Related Works
Studies utilizing textual prompting with direct integration of time-series data represent the numerical data in a structured text format for the model. One of the earliest studies in this context [5] demonstrated the use of LLMs for tasks such as arrhythmia detection and physical activity recognition by directly integrating interbeat interval sequences and accelerometer sensor data into prompts. Similarly, Kim et al. [6] explored LLMs for health prediction by incorporating various sensor readings into queries, such as step count, calories burned, and resting heart rate. Chan et al. [4] proposed an LLM-based framework for tasks such as semantic segmentation, boundary detection, and anomaly detection, embedding biosignals directly into model inputs. Additionally, Imran et al. [7] employed LLMs for physical activity analysis, where gyroscope and accelerometer sensor data were incorporated into prompts. LLMs have also been applied to cardiac activity monitoring, metabolic health prediction, and sleep detection by leveraging wearable sensor data [8]. However, LLMs are not inherently specialized for processing long raw numerical sequences. Physiological time-series data often consist of thousands or even millions of data points, and representing them as textual input for LLMs can result in poor performance. This approach also exceeds the token limits of most LLMs and leads to high computational costs due to the large number of tokens. For instance, the input token limit of GPT-4 models is 128,000 tokens, while a 16-minute PPG signal sampled at just 20 Hz results in over 130,000 tokens.
@@ III-A2 Orchestrator
The Orchestrator handles failures (e.g., in data retrieval or task execution) through iterative communication between the Task Planner and Task Executor. When a step fails, the orchestrator can dynamically replan by revisiting earlier stages or prompting the user for clarification. Additionally, the agent is not limited to specific predefined tasks. OpenCHA’s modular design enables the orchestrator to invoke any domain-specific tools based on the user’s query. This design ensures flexibility and extensibility to handle a broader range of user queries by simply registering additional analytical tools as new tasks within the framework.
@@ IV Evaluation and Results
We evaluate our LLM-powered agent using the dataset described in Section III-B1. For this evaluation, we use PPG data from 12 participants, comprising 621 PPG recordings, each spanning 15 minutes. These 12 participants were held out as a test set and were not used in any stage of model development or fine-tuning in the PPG processing pipeline. The remaining 34 participants’ data in the dataset were used exclusively for training and validation of the machine learning models in the pipeline. The pipeline was fully finalized before evaluating the 12 held-out participants to ensure an unbiased assessment of the agent’s performance.
@@ IV-B Gold reference and comparison criteria
For evaluation, we compute Mean Absolute Error (${MAE}=\frac{1}{n}\sum_{i=1}^{n}|y_{i}-\hat{y}_{i}|$), Root Mean Square Error (${RMSE}=\sqrt{\frac{1}{n}\sum_{i=1}^{n}(y_{i}-\hat{y}_{i})^{2}}$), Mean Absolute Percentage Error (${MAPE}=\frac{100\%}{n}\sum_{i=1}^{n}\left|\frac{y_{i}-\hat{y}_{i}}{y_{i}}\right|$), and Median Absolute Difference (${MAD}=\text{median}(|y_{i}-\hat{y}_{i}|)$), where $n$ is the number of samples, $y_{i}$ represents the ground-truth values, and $\hat{y}_{i}$ is the predicted value of HR. Additionally, we perform linear regression analysis to assess the linear relationship between estimated and reference HR values and use the Bland-Altman method [30] to evaluate the agreement between the estimated and ground truth HR values.
@@ IV-C Results and Performance Analysis
The higher outlier rates in the OpenAI model approaches are due to the inherent uncertainty in LLM-generated and executed codes. Although Assistants execute codes in a controlled environment, the accuracy of the analysis depends on the correctness of the code generated. In many cases, the OpenAI models produced incomplete or inaccurate signal processing codes, which led to outliers in HR estimation. In contrast, our orchestrated agent employs validated and reliable processing tools, ensuring robust analysis and physiologically realistic results.